# Joint Transmission and Compression Design for 6G Networked Sensing with Limited-Capacity Backhaul

### v1

@@ Abstract
This paper considers networked sensing in cellular network, where multiple base stations (BSs) first compress their received echo signals from multiple targets and then forward the quantized signals to the cloud via limited-capacity backhaul links, such that the cloud can leverage all useful echo signals to perform high-resolution localization. Under this setup, we manage to characterize the posterior Cramér-Rao Bound (PCRB) for localizing all the targets as a function of the transmit covariance matrix and the compression noise covariance matrix of each BS. Then, a PCRB minimization problem subject to the transmit power constraints and the backhaul capacity constraints is formulated to jointly design the BSs’ transmission and compression strategies. We propose an efficient algorithm to solve this problem based on the alternating optimization technique. Specifically, it is shown that when either the transmit covariance matrices or the compression noise covariance matrices are fixed, the successive convex approximation technique can be leveraged to optimize the other type of covariance matrices locally. Numerical results are provided to verify the effectiveness of our proposed algorithm.
@@ I Introduction
Integrated sensing and communication (ISAC) is identified as one of the six key usage scenarios of the six-generation (6G) cellular systems [1]. The main challenge in ISAC is how to embed the sensing functionality into the future 6G communication systems efficiently [2, 3]. In the literature, plenty of works have studied the practical strategies to perform sensing[4, 5, 6] and the fundamental limits of sensing [7, 8, 9, 10, 11] in communication systems. It is worth noting that in most of the existing ISAC works, the sensing functionality is achieved via one base station (BS), just as in radar systems.
@@
However, there exists a notable difference between the BSs and the conventional radars - the BSs across large-scale area are inter-connected via the backhaul links and can share information with each other. Actually, this property has been utilized to achieve cooperative communication in cellular network, giving rise to cloud radio access network (C-RAN), cell-free massive multiple-input multiple-output (MIMO), etc., where multiple BSs jointly serve the users via sharing the data messages and the channel state information (CSI). Similarly, in the future 6G ISAC systems, multiple BSs may share the local sensing information with each other to perform networked sensing [12, 13]. This is a major advantage of 6G sensing systems over the conventional radar systems.
In this paper, we consider a 6G networked sensing system consisting of multiple BSs and multiple targets, as shown in Fig. 1. Each BS first emits wireless signals in the downlink and then receives the echo signals from the targets. Next, each BS adopts a compress-then-forward strategy, under which it first compresses its received echo signals and then forwards the compression bits to the cloud via a limited-capacity backhaul link. At last, the cloud may utilize the global information observed by all the BSs to jointly localize the targets. Under this system, we manage to characterize the Posterior Cramér-Rao Bound (PCRB) [14] for localization as a function of the BSs’ transmit covariance matrices and compression noise covariance matrices. Based on this result, we formulate a PCRB minimization problem subject to each BS’s transmit power constraint and backhaul capacity constraint. Moreover, we propose an efficient algorithm to solve this problem via the alternating optimization technique. Numerical results are provided to verify the effectiveness of our proposed algorithm. In the literature, most of the ISAC works focus on the sensing limits in a single-BS and single-target scenario [7, 8, 9, 10, 11], where sensing is performed locally by the BS. Our work makes an early attempt to consider PCRB characterization under multi-BS and multi-target scenario and investigate the impact of limited-capacity backhaul on sensing design.
@@ II System Model
Figure 1: System model for networked sensing with limited-capacity backhaul.
We consider the sensing job in a multi-cell ISAC system as shown in Fig. 1, where $N$ BSs, each equipped with $M_{t}$ transmit antennas and $M_{r}$ receive antennas, are connected to the cloud via the backhaul links and cooperatively localize $K$ targets via utilizing the wireless signals. We focus on a two-dimensional (2D) coordinate system, while the known position of BS $n$ is denoted as $\bm{\rho}_{n}=[\rho_{n}^{x},\rho_{n}^{y}]^{T}\in\mathbb{R}^{2\times 1}$, and the unknown position of target $k$ is denoted as $\bm{q}_{k}=[q_{k}^{x},q_{k}^{y}]^{T}\in\mathbb{R}^{2\times 1}$, $n=1,\ldots,N$ and $k=1,\ldots,K$. In this paper, we assume that the unknown target positions $\bm{q}=[\bm{q}_{1}^{T},\dots,\bm{q}_{K}^{T}]^{T}\in\mathbb{R}^{2K\times 1}$ are random with a joint probability density function (PDF) of $p_{\bm{q}}(\bm{q})$. Moreover, the angle-of-departure (AOD) of the signal transmitted from BS $n$ to target $k$ and the angle-of-arrival (AOA) of the signal reflected from target $k$ to BS $n$ are denoted as
@@
Moreover, define $\bm{v}(\theta_{n,k})\in\mathbb{C}^{M_{t}\times 1}$ and $\bm{a}(\theta_{n,k})\in\mathbb{C}^{M_{r}\times 1}$ as the transmit and receive steering vectors of BS $n$ towards target $k$, $\forall n,k$.
Under the above system, all the BSs transmit wireless signals simultaneously in the downlink, then receive the echo signals from the targets, and forward their received echo signals to the cloud via their backhaul links for performing localization. Specifically, let $\bm{x}_{n}\in\mathbb{C}^{M_{t}\times 1}$ with the covariance matrix $\bm{R}_{n}=\mathbb{E}[\bm{x}_{n}\bm{x}_{n}^{H}]$ denote the signal emitted by the transmit antenna array of BS $n$, $\forall n$. Then, the echo signals at the receive antenna array of BS $n$ can be written as
$\bm{y}_{n}\!=\!\bm{A}_{n}\sum_{u=1}^{N}\bm{B}_{n,u}\bm{V}_{u}^{T}\bm{x}_{u}+\bm{z}_{n}\!=\!\sum_{u=1}^{N}\bm{G}_{n,u}\bm{x}_{u}+\bm{z}_{n},\forall n.$ (2)
In the above, $\bm{V}_{n}=[\bm{v}(\theta_{n,1}),\ldots,\bm{v}(\theta_{n,K})]\in\mathbb{C}^{M_{t}\times K}$ and $\bm{A}_{n}=[\bm{a}(\theta_{n,1}),\ldots,\bm{a}(\theta_{n,K})]\in\mathbb{C}^{M_{r}\times K}$ denote the collection of the transmit and receive steering vectors of BS $n$, $\bm{B}_{n,u}=\textrm{diag}(\bm{b}_{n,u})\in\mathbb{C}^{K\times K}$, where $\bm{b}_{n,u}=[b_{n,u,1},\ldots,b_{n,u,K}]^{T}$ with $b_{n,u,k}$ being the attenuation coefficient from the transmit antenna array of BS $u$ to target $k$ to the receive antenna array of BS $n$, which contains both the round-trip path loss and the radar cross-section (RCS) of the target, $\bm{z}_{n}\sim\mathcal{CN}(0,\sigma^{2}\bm{I})$ is the Gaussian noise at BS $n$, and $\bm{G}_{n,u}\triangleq\bm{A}_{n}\bm{B}_{n,u}\bm{V}_{u}^{T}$, $\forall n,u$.
Due to the limited backhaul capacity, BS $n$ needs to compress the echo signal $\bm{y}_{n}$, $\forall n$. In this paper, we consider the Gaussian test channel to model the compression process [15]. Specifically, the quantized signal of BS $n$ is given as [15]
@@
where $\bm{e}_{n}\in\mathbb{C}^{M_{r}\times 1}$ is the Gaussian quantization error vector with zero mean and covariance matrix $\bm{Q}_{n}\in\mathbb{C}^{M_{r}\times M_{r}}$, i.e., $\bm{e}_{n}\sim\mathcal{CN}(\bm{0},\bm{Q}_{n})$. Based on (3), the rate (in terms of bits/second/sample) for BS $n$ to transmit $\tilde{\bm{y}}_{n}$ via backhaul is
@@
where the expectation is over the target location vector $\bm{q}=[\bm{q}_{1}^{T},\dots,\bm{q}_{K}^{T}]^{T}$, which determines the channel $\bm{G}_{n,u}$’s.
After receiving the quantized signals from all the $N$ BSs, the cloud aims to localize all the $K$ targets. In the following, we first characterize the localization performance under networked sensing as a function of $\bm{R}_{1},\ldots,\bm{R}_{N}$ and $\bm{Q}_{1},\ldots,\bm{Q}_{N}$, and then optimize these covariance matrices to improve the sensing performance subject to the backhaul capacity constraints.
@@ III Posterior Cramér-Rao Bound for Localization
In this work, we utilize the PCRB [14] as the performance metric to evaluate the performance of the networked sensing system described in the above. Specifically, let $\bm{q}=[\bm{q}_{1}^{T},\ldots,\bm{q}_{K}^{T}]^{T}$ denote the collection of target locations, and $\bm{b}^{\rm R}=[(\bm{b}^{\rm R}_{1,1})^{T},\ldots,(\bm{b}^{\rm R}_{1,N})^{T},\ldots,(\bm{b}^{\rm R}_{N,1})^{T},\ldots,(\bm{b}^{\rm R}_{N,N})^{T}]^{T}$ and $\bm{b}^{\rm I}=[(\bm{b}^{\rm I}_{1,1})^{T},\ldots,(\bm{b}^{\rm I}_{1,N})^{T},\ldots,(\bm{b}^{\rm I}_{N,1})^{T},\ldots,(\bm{b}^{\rm I}_{N,N})^{T}]^{T}$ denote the collection of the real parts and that of the imaginary parts of the attenuation coefficients, where $\bm{b}^{\rm R}_{n,u}=[b_{n,,1}^{\rm R},\ldots,b_{n,u,K}^{\rm R}]^{T}$ and $\bm{b}^{\rm I}_{n,u}=[b_{n,u,1}^{\rm I},\ldots,b_{n,u,K}^{\rm I}]^{T}$ with $b_{n,u,k}^{\rm R}$ and $b_{n,u,k}^{\rm I}$ being the real part and the imaginary part of $b_{n,u,k}$, $\forall n,u,k$. Moreover, define $\boldsymbol{\xi}=[\bm{q}^{T},(\bm{b}^{\rm R})^{T},(\bm{b}^{\rm I})^{T}]^{T}$ as the collection of all the unknown parameters to be estimated, and $\tilde{\bm{y}}=[\tilde{\bm{y}}_{1}^{T},\ldots,\tilde{\bm{y}}_{N}^{T}]^{T}$ as the collection of the de-quantized signals at the cloud, where $\tilde{\bm{y}}_{n}$ is given in (3). We assume that the target locations are random variables with a joint probability density function (PDF) $p_{\bm{q}}(\bm{q})$, and the attenuation coefficients are deterministic but unknown variables. In the rest of this section, we aim to characterize the PCRB for estimating $\boldsymbol{\xi}$ based on the de-quantized signals at the cloud $\tilde{\bm{y}}$ and the prior knowledge of PDF $p_{\bm{q}}(\bm{q})$.
In order to derive the PCRB for $\boldsymbol{\xi}$, we first calculate the Fisher information matrix (FIM) for estimating $\boldsymbol{\xi}$
$\bm{F}_{\boldsymbol{\xi}}(\{\bm{R}_{n},\bm{Q}_{n}\}_{n=1}^{N})=\bm{F}_{0,\boldsymbol{\xi}}+\bm{F}_{p},$ (5)
where $\bm{F}_{0,\boldsymbol{\xi}}$ denotes the FIM corresponding to the de-quantized signals given in (3)
$\bm{F}_{0,\boldsymbol{\xi}}=\mathbb{E}_{\tilde{\bm{y}},\bm{q}}\left[\frac{\partial\log p(\tilde{\bm{y}}|\boldsymbol{\xi})}{\partial\boldsymbol{\xi}}\Bigg{(}\frac{\partial\log p(\tilde{\bm{y}}|\boldsymbol{\xi})}{\partial\boldsymbol{\xi}}\Bigg{)}^{T}\right],$ (6)
@@
$p(\tilde{\bm{y}}|\boldsymbol{\xi})=$ $~{}\prod_{n=1}^{N}\frac{1}{\pi^{L}|\sigma^{2}\bm{I}+\bm{Q}_{n}|}\exp\Bigg{[}-\Bigg{(}\tilde{\bm{y}}_{n}-\sum_{u=1}^{N}\bm{G}_{n,u}\bm{x}_{u}\Bigg{)}^{H}$
@@
$\bm{F}_{p}$ $=\mathbb{E}_{\bm{q}}\left[\frac{\partial\log p_{\bm{q}}(\bm{q})}{\partial\boldsymbol{\xi}}\Bigg{(}\frac{\partial\log p_{\bm{q}}(\bm{q})}{\partial\boldsymbol{\xi}}\Bigg{)}^{T}\right]$
@@
Given any PDF $p_{\bm{q}}(\bm{q})$, $\bm{F}_{p}$ can be computed based on (III). Therefore, the remaining job to characterize $\bm{F}_{\boldsymbol{\xi}}(\{\bm{R}_{n},\bm{Q}_{n}\}_{n=1}^{N})$ is to quantify $\bm{F}_{0,\boldsymbol{\xi}}$ given in (6). It is observed from (2) that $\bm{G}_{n,u}$’s are functions of $\bm{\theta}=[\theta_{1,1},\ldots,\theta_{1,N},\ldots,\theta_{N,1},\ldots,\theta_{N,K}]^{T}$. Moreover, because $\bm{\theta}$ is a complicated function of the target locations $\bm{q}$ as shown in (1), $p(\tilde{\bm{y}}|\boldsymbol{\xi})$ is a complicated function of $\bm{q}$, and it is challenging to calculate its derivative over $\bm{q}$. We apply the chain rule to tackle this challenge. Specifically, define $\boldsymbol{\zeta}=[\boldsymbol{\theta}^{T},(\bm{b}^{\rm R})^{T},(\bm{b}^{\rm I})^{T}]^{T}$. Moreover, define
$\bm{u}_{n,k}=\left[\frac{\partial\theta_{n,k}}{\partial q_{k}^{x}},\frac{\partial\theta_{n,k}}{\partial q_{k}^{y}}\right]^{T},~{}\forall n,k,$ (9)
@@
Then, based on the chain rule, $\bm{F}_{0,\boldsymbol{\xi}}$ can be obtained as follows
$\bm{F}_{0,\boldsymbol{\xi}}=\mathbb{E}_{\boldsymbol{\xi}}[\bm{U}\bm{F}_{0,\boldsymbol{\zeta}}\bm{U}^{T}].$ (12)

### v3

# Joint Transmission and Compression Optimization for Networked Sensing with Limited-Capacity Fronthaul Links

@@ Abstract
This paper considers networked sensing in cellular network, where multiple base stations (BSs) first compress their received echo signals from multiple targets and then forward the quantized signals to the central unit (CU) via limited-capacity fronthaul links, such that the CU can leverage all useful echo signals to perform high-resolution localization. Under this setup, we manage to characterize the posterior Cramér-Rao Bound (PCRB) for localizing all the targets with random positions as a function of the transmit covariance matrix and the compression noise covariance matrix of each BS. Then, a PCRB minimization problem subject to the transmit power constraints and the fronthaul capacity constraints is formulated to jointly design the BSs’ transmission and compression strategies. We propose an efficient algorithm to solve this problem based on the alternating optimization technique. Specifically, it is shown that when either the transmit covariance matrices or the compression noise covariance matrices are fixed, the successive convex approximation (SCA) technique can be leveraged to optimize the other type of covariance matrices locally optimally. Moreover, we also propose a novel estimate-then-beamform-then-compress strategy for the massive receive antenna scenario, under which each BS first estimates targets’ angle-of-arrivals (AOAs) locally, then beamforms its high-dimension received signals into low-dimension ones based on the estimated AOAs, and last compresses the beamformed signals for fronthaul transmission. An efficient beamforming and compression design method is devised under this strategy. Numerical results are provided to verify the effectiveness of our proposed algorithms.
Index Terms:
Networked sensing, integrated sensing and communication (ISAC), posterior Cramér-Rao Bound (PCRB), limited-capacity fronthaul, alternating optimization.
@@ I Introduction
Integrated sensing and communication (ISAC) is identified as one of the six key usage scenarios of the six-generation (6G) cellular systems [2]. The main challenge in ISAC is how to embed the sensing functionality into the future 6G communication systems efficiently [3, 4]. In the literature, plenty of works have studied the practical strategies to perform sensing [5, 6, 7] and the fundamental limits of sensing [8, 9, 10] in communication systems. It is worth noting that in most of the existing ISAC works, the sensing functionality is achieved via one base station (BS), just as in radar systems.
@@
However, there exists a notable difference between the BSs and the conventional radars - the BSs across large-scale area are inter-connected via the fronthaul links and can share information with each other. Actually, this property has been utilized to achieve cooperative communication in cellular network, giving rise to cloud radio access network (C-RAN), cell-free massive multiple-input multiple-output (MIMO), etc., where multiple BSs jointly serve the users via sharing the data messages and the channel state information (CSI). Similarly, in the future 6G ISAC systems, multiple BSs may share the local sensing information with each other to perform networked sensing [11]. This is a major advantage of 6G sensing systems over the conventional radar systems.
Motivated by the cooperation gain in communication, this paper considers the networked sensing technique, under which multiple BSs forward their received echo signals reflected by multiple targets to a central unit (CU), and the CU jointly localizes all the targets based on the global echo signals. Theoretically speaking, the sensing performance can be significantly enhanced by fusing the global sensing information at the CU. However, in practice, the BSs are connected to the CU via limited-capacity fronthaul links and they can only transmit compression bits over these links. Therefore, the echo signals received by the CU are subject to compression noise, whose power depends on the compression strategy adopted by each BS. This indicates that besides the transmission strategy that is considered in single-BS based sensing scenario, each BS should also devise its compression strategy carefully under networked sensing. In this paper, we make an early attempt to jointly design the transmission and compression strategies of each BS under networked sensing with limited-capacity fronthaul.
I-A Related Works
The functionality of sensing is already recognized to be indispensable in the future 6G wireless systems and has attracted many interests of researchers. Under the single-BS sensing setup, previous works have investigated the fundamental limits of sensing [12, 13, 14, 15, 16, 17, 18, 19, 20, 21] to provide useful insights for the practical sensing technique design. Therein, the works [12, 13, 14, 15, 16, 17] derive the Cramér-Rao Bound (CRB) to characterize the localization performance with unknown but deterministic target locations. Then, the works [18, 20, 19, 21] consider the scenario where the target locations are random with known distribution and give the explicit expression of the posterior Cramér-Rao Bound (PCRB) on target directions by exploiting the location distribution. By regarding the CRB or the PCRB as the objective function, the works [13, 14, 15, 16, 17, 18, 19, 21, 20] propose to optimize the waveform to realize the optimal sensing performance under certain constraints. Besides the single-BS setup, the performance limits in terms of the CRB for target localization under networked sensing are also studied in [22, 23]. Specifically, the work [22] considers the system where both the transmitters and receivers are equipped with one single antenna and analyzes the CRB for wideband localization. Then, the work [23] derives the CRB for the cooperative ISAC system where multiple BSs localizes the targets by estimating their directions and goes forward to optimize the transmit beamformers to balance the trade-off between sensing CRB and transmission rate. However, both of these two works do not take the practical limited-capacity fronthaul issue in cooperative networks [24, 25, 26] into consideration. In networked sensing systems, the transmit strategy and the compression strategy of the BSs should be jointly optimized.
Several works make efforts in studying the fundamentals of networked sensing by considering the effect of compression arising from fronthaul links. The work [27] adopts the information-theoretic criterion of the Bhattacharyya distance and quantization rate to measure the detection performance and backhaul capacity, respectively, in the cloud radio-multistatic radar system, which are then used for the joint optimization of the temporal code vector and backhaul quantization. Furthermore, the work [28] derives the expected CRB for cooperative localization and considers the optimization of the bit allocation among the receive BSs. However, the works [27, 28] consider the case of one transmit BS and multiple receive BSs, with each BS equipped with one antenna. In practical multi-cell communication systems, multiple BSs will transmit at the same time, resulting in inter-cell echo signals for localization. Such signals can make the characterization of localization PCRB quite challenging, which is not considered in the above works. Moreover, current BSs are equipped with multiple antennas. On one hand, targets’ angle information can be extracted with multi-antenna receive BSs, which is not considered in [27, 28]. On the other hand, due to the multiple antennas at the BSs, the transmit beamforming strategy should be jointly optimized with the receive compression strategy, which is not considered in [27, 28].
Figure 1: System model for networked sensing with limited-capacity fronthaul.
I-B Main Contributions
In this paper, we consider a 6G networked sensing system consisting of multiple BSs, each equipped with multiple antennas, and multiple targets, as shown in Fig. 1. Each BS first emits wireless signals in the downlink and then receives the echo signals from the targets. Next, each BS adopts a compress-then-forward strategy, under which it first compresses its received echo signals and then forwards the compression bits to the CU via a limited-capacity fronthaul link. At last, the CU may utilize the global information observed by all the BSs to jointly localize the targets. Our main contributions are summarized as follows:
  • Under the networked sensing setup with multiple BSs equipped with multiple antennas and multiple randomly located targets, we manage to characterize the PRCB for localization as a function of the BSs’ transmit covariance matrices and compression noise covariance matrices. Such a result can be viewed as a generalization of the classic result under [13] which characterizes the CRB with a single anchor.
  • Next, we formulate a joint transmission and compression optimization problem, which minimizes the PCRB subject to each BS’s transmit power constraint and fronthaul capacity constraint. The formulated problem is non-convex and quite challenging to solve, due to the coupling of the transmit covariance matrices and compression noise covariance matrices. To deal with the these issues, we first propose an alternating optimization (AO)-based method to separately optimize the transmit covariance matrices and the compression noise covariance matrices. Moreover, we show that when each type of covariance matrices is fixed, we can apply the successive convex approximation (SCA) technique to optimize the other type locally optimally.
  • We also propose an estimate-then-beamform-then-compress (EBC) strategy to reduce the algorithm complexity when the number of receive antennas is large. Specifically, under the proposed EBC strategy, each BS first estimates the angle-of-arrivals (AOAs) of the targets, and then exploits the AOA information to beamform the high-dimension received signals onto reduced dimensions, each corresponding to the main AOA of some target, and last compresses the lower-dimension beamformed signals for fronthaul transmission. This method is shown numerically to achieve very attractive performance with significantly reduced computational complexity.
I-C Organizations and Notations
The rest of the paper is organized as follows. Section II gives the introduction of the system model and Section III derives the expression of the PCRB for localization in the considered system. In Section IV, the problem of joint transmission and compression optimization is formulated and the AO-based algorithm is proposed to solve it. Then an EBC strategy is introduced in Section V. Finally, Section VI evaluates the performance of the proposed algorithms and Section VII concludes this paper.
In this paper, vectors and matrices are denoted by boldface lower-case letters and boldface upper-case letters, respectively. For a complex vector $\bm{x}$, $||\bm{x}||_{q}$ and $x_{n}$ denote the $l_{q}$-norm and the $n$-th element, respectively. For an $M\times N$ matrix $\bm{X}$, $\bm{X}^{T}$, $\bm{X}^{*}$, $\bm{X}^{H}$, and $\bm{X}(m,n)$ denote its transpose, conjugate, conjugate transpose and element located at the $m$th row and $n$th column, respectively. For a square matrix $\bm{A}$, $|\bm{A}|$, $\text{tr}(\bm{A})$, and $\bm{A}^{-1}$ denote its determinant, trace, and inverse, respectively, and $\bm{A}\succeq\bm{0}$ means that $\bm{A}$ is positive semidefinite. The distribution of a circularly symmetric complex Gaussian random vector with mean $\bm{\mu}$ and covariance $\bm{\Sigma}$ is denoted by $\mathcal{CN}(\bm{\mu},\bm{\Sigma})$. Further, $\mathbb{E}_{\bm{a}}[\cdot]$ denotes the expectation operation over random vector $\bm{a}$, operation $\odot$ denotes Hadamard product of two matrices, and ${}\mathcal{R}\{\cdot\}$ denote the column space of the argument. Letters $\mathbb{R}^{M\times N}$ and $\mathbb{C}^{M\times N}$ denote the region of real matrix and complex matrix, respectively, which have the size of $M\times N$. Finally, letters $\mathbb{S}^{N}_{+}$ and $\mathbb{S}^{N}_{++}$ denote the region of positive semidefinite matrices and positive definite matrices, respectively, with size $N\times N$.
@@ II System Model
We consider the sensing job in a multi-cell ISAC system as shown in Fig. 1, where $N$ BSs, each equipped with $M_{t}$ transmit antennas and $M_{r}$ receive antennas, are connected to the CU via the fronthaul links and cooperatively localize $K$ targets via utilizing the wireless signals. We focus on a two-dimensional (2D) coordinate system, while the known position of BS $n$ is denoted as $\bm{\rho}_{n}=[\rho_{n}^{x},\rho_{n}^{y}]^{T}\in\mathbb{R}^{2\times 1}$, and the unknown position of target $k$ is denoted as $\bm{q}_{k}=[q_{k}^{x},q_{k}^{y}]^{T}\in\mathbb{R}^{2\times 1}$, $n=1,\ldots,N$ and $k=1,\ldots,K$. In this paper, we assume that the unknown target positions $\bm{q}=[\bm{q}_{1}^{T},\dots,\bm{q}_{K}^{T}]^{T}\in\mathbb{R}^{2K\times 1}$ are random with a joint probability density function (PDF) of $p_{\bm{q}}(\bm{q})$. Moreover, the angle-of-departure (AOD) of the signal transmitted from BS $n$ to target $k$ and the angle-of-arrival (AOA) of the signal reflected from target $k$ to BS $n$ are denoted as
@@
For convenience, in the rest of this paper, we define $\theta_{n,k}$ as the angle between BS $n$ and target $k$. Moreover, define $\bm{v}(\theta_{n,k})\in\mathbb{C}^{M_{t}\times 1}$ and $\bm{a}(\theta_{n,k})\in\mathbb{C}^{M_{r}\times 1}$ as the transmit and receive steering vectors of BS $n$ towards target $k$, $\forall n,k$.
Under the above system, all the BSs transmit wireless signals simultaneously in the downlink, then receive the echo signals from the targets, and forward their received echo signals to the CU via their fronthaul links for performing localization. Specifically, let $\bm{x}_{n}\in\mathbb{C}^{M_{t}\times 1}$ with the covariance matrix $\bm{R}_{n}=\mathbb{E}[\bm{x}_{n}\bm{x}_{n}^{H}]$ denote the signal emitted by the transmit antenna array of BS $n$, $\forall n$. Then, the echo signals at the receive antenna array of BS $n$ can be written as
$\bm{y}_{n}$ $=\bm{A}_{n}\sum_{u=1}^{N}\bm{B}_{n,u}\bm{V}_{u}^{T}\bm{x}_{u}+\bm{z}_{n}$
$=\sum_{u=1}^{N}\bm{G}_{n,u}\bm{x}_{u}+\bm{z}_{n},~{}\forall n.$ (2)
In the above, $\bm{V}_{n}=[\bm{v}(\theta_{n,1}),\ldots,\bm{v}(\theta_{n,K})]\in\mathbb{C}^{M_{t}\times K}$ and $\bm{A}_{n}=[\bm{a}(\theta_{n,1}),\ldots,\bm{a}(\theta_{n,K})]\in\mathbb{C}^{M_{r}\times K}$ denote the collection of the transmit and receive steering vectors of BS $n$, $\bm{B}_{n,u}=\textrm{diag}(\bm{b}_{n,u})\in\mathbb{C}^{K\times K}$, where $\bm{b}_{n,u}=[b_{n,u,1},\ldots,b_{n,u,K}]^{T}\in\mathbb{C}^{K\times 1}$ with $b_{n,u,k}$ being the unknown but deterministic attenuation coefficient from the transmit antenna array of BS $u$ to target $k$ to the receive antenna array of BS $n$ containing both the round-trip path loss and the radar cross-section (RCS) of the target, $\bm{z}_{n}\sim\mathcal{CN}(0,\sigma^{2}\bm{I})$ is the Gaussian noise at BS $n$, and
$\bm{G}_{n,u}\triangleq\bm{A}_{n}\bm{B}_{n,u}\bm{V}_{u}^{T},~{}\forall n,u.$ (3)
Due to the limited fronthaul capacity, each BS $n$ needs to compress the echo signal $\bm{y}_{n}$ and send the quantization bits to the CU over the fronthaul link, $\forall n$. In this paper, we consider the Gaussian test channel to model the compression process [29]. Specifically, the quantized signal of BS $n$ is given as
@@
where $\bm{e}_{n}\in\mathbb{C}^{M_{r}\times 1}$ is the Gaussian quantization error vector with zero mean and covariance matrix $\bm{Q}_{n}\in\mathbb{C}^{M_{r}\times M_{r}}$, i.e., $\bm{e}_{n}\sim\mathcal{CN}(\bm{0},\bm{Q}_{n})$. Based on (4), the rate (in terms of bits per second per sample) for BS $n$ to transmit $\tilde{\bm{y}}_{n}$ to the CU via the fronthaul link is
@@
where the expectation is over the target location vector $\bm{q}$, which determines the channel $\bm{G}_{n,u}$’s.
After receiving the quantized signals from all the $N$ BSs, the CU aims to localize all the $K$ targets. In the following, we first characterize the localization performance under the above networked sensing system as a function of the transmit covariance matrices $\bm{R}_{1},\ldots,\bm{R}_{N}$ and the compression noise covariance matrices $\bm{Q}_{1},\ldots,\bm{Q}_{N}$, and then optimize the transmit and compression noise covariance matrices to improve the sensing performance subject to the fronthaul capacity constraints.
@@ III Posterior Cramér-Rao Bound for Localization
In this work, we utilize the PCRB as the performance metric to evaluate the performance of the networked sensing system described in the above [30]. Specifically, let $\bm{b}^{\rm R}=[(\bm{b}^{\rm R}_{1,1})^{T},\ldots,(\bm{b}^{\rm R}_{1,N})^{T},\ldots,(\bm{b}^{\rm R}_{N,1})^{T},\ldots,(\bm{b}^{\rm R}_{N,N})^{T}]^{T}$ and $\bm{b}^{\rm I}=[(\bm{b}^{\rm I}_{1,1})^{T},\ldots,(\bm{b}^{\rm I}_{1,N})^{T},\ldots,(\bm{b}^{\rm I}_{N,1})^{T},\ldots,(\bm{b}^{\rm I}_{N,N})^{T}]^{T}$ denote the collection of the real parts and that of the imaginary parts of the attenuation coefficients, respectively, where $\bm{b}^{\rm R}_{n,u}=[b_{n,,1}^{\rm R},\ldots,b_{n,u,K}^{\rm R}]^{T}$ and $\bm{b}^{\rm I}_{n,u}=[b_{n,u,1}^{\rm I},\ldots,b_{n,u,K}^{\rm I}]^{T}$ with $b_{n,u,k}^{\rm R}$ and $b_{n,u,k}^{\rm I}$ being the real part and the imaginary part of $b_{n,u,k}$, $\forall n,u,k$. Moreover, define $\bm{\xi}=[\bm{q}^{T},(\bm{b}^{\rm R})^{T},(\bm{b}^{\rm I})^{T}]^{T}$ as the collection of all the unknown parameters to be estimated, and $\tilde{\bm{y}}=[\tilde{\bm{y}}_{1}^{T},\ldots,\tilde{\bm{y}}_{N}^{T}]^{T}$ as the collection of the de-quantized signals at the CU, where $\tilde{\bm{y}}_{n}$ is given in (4). In the rest of this section, we aim to characterize the PCRB for estimating $\bm{\xi}$ based on the de-quantized signals at the CU $\tilde{\bm{y}}$ and the prior knowledge of PDF $p_{\bm{q}}(\bm{q})$.
In order to derive the PCRB for $\bm{\xi}$, we first calculate the posterior Fisher information matrix (PFIM) for estimating $\bm{\xi}$
$\bm{F}_{\bm{\xi}}(\{\bm{R}_{n},\bm{Q}_{n}\}_{n=1}^{N})=\bm{F}_{0,\bm{\xi}}+\bm{F}_{p},$ (6)
where $\bm{F}_{0,\bm{\xi}}$ denotes the FIM corresponding to the de-quantized signals given in (4)
$\bm{F}_{0,\bm{\xi}}=\mathbb{E}_{\tilde{\bm{y}},\bm{q}}\left[\frac{\partial\log p(\tilde{\bm{y}}|\bm{\xi})}{\partial\bm{\xi}}\Bigg{(}\frac{\partial\log p(\tilde{\bm{y}}|\bm{\xi})}{\partial\bm{\xi}}\Bigg{)}^{T}\right],$ (7)
@@
$p(\tilde{\bm{y}}|\bm{\xi})=$ $~{}\prod_{n=1}^{N}\frac{1}{\pi^{M_{r}}|\sigma^{2}\bm{I}+\bm{Q}_{n}|}\exp\Bigg{[}-\Bigg{(}\tilde{\bm{y}}_{n}-\sum_{u=1}^{N}\bm{G}_{n,u}\bm{x}_{u}\Bigg{)}^{H}$
@@
$\bm{F}_{p}$ $=\mathbb{E}_{\bm{q}}\left[\frac{\partial\log p_{\bm{q}}(\bm{q})}{\partial\bm{\xi}}\Bigg{(}\frac{\partial\log p_{\bm{q}}(\bm{q})}{\partial\bm{\xi}}\Bigg{)}^{T}\right]$
@@
Given any PDF $p_{\bm{q}}(\bm{q})$, $\bm{F}_{p}$ can be computed based on (III). Therefore, the remaining job to characterize $\bm{F}_{\bm{\xi}}(\{\bm{R}_{n},\bm{Q}_{n}\}_{n=1}^{N})$ is to quantify $\bm{F}_{0,\bm{\xi}}$ given in (7). It is observed from (3) that $\bm{G}_{n,u}$’s are functions of $\bm{\theta}=[\bm{\theta}_{1}^{T},\dots,\bm{\theta}_{N}^{T}]^{T}$ with $\bm{\theta}_{n}=[\theta_{n,1},\ldots,\theta_{n,K}]^{T}$. Moreover, because $\bm{\theta}$ is a complicated function of the target locations $\bm{q}$ as shown in (1), $p(\tilde{\bm{y}}|\bm{\xi})$ is a complicated function of $\bm{q}$, and it is challenging to calculate its derivative over $\bm{q}$. We apply the chain rule to tackle this challenge. Specifically, define $\bm{\zeta}=[\bm{\theta}^{T},(\bm{b}^{\rm R})^{T},(\bm{b}^{\rm I})^{T}]^{T}$. Note that different from parameter vector $\bm{\xi}$, target location $\bm{q}$ has been changed to target angle $\bm{\theta}$ in the parameter vector $\bm{\zeta}$. Moreover, define
$\bm{u}_{n,k}=\frac{\partial\theta_{n,k}}{\partial\bm{q}_{k}}=\left[\frac{\partial\theta_{n,k}}{\partial q_{k}^{x}},\frac{\partial\theta_{n,k}}{\partial q_{k}^{y}}\right]^{T},~{}\forall n,k,$ (10)
@@
Then, based on the chain rule, $\bm{F}_{0,\bm{\xi}}$ can be obtained as follows
$\bm{F}_{0,\bm{\xi}}=\mathbb{E}_{\bm{\xi}}[\bm{U}\bm{F}_{0,\bm{\zeta}}\bm{U}^{T}].$ (13)